OpenStudy (anonymous):

Factor 3x^9+6x^8-9x^7+6x^6

OpenStudy (anonymous):

http://mathway.com/ that website is good for things like this. it gives you different options on how to do it and everything.

OpenStudy (anonymous):

Simplify 3x9 + 6x8 + -9x7 + 6x6 = 0 Reorder terms 6x6 + -9x7 + 6x8 + 3x9 = 0 Solve 6x6 + -9x7 + 6x8 + 3x9 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '3x6'. 3x6(2 + -3x + 2x2 + x3) = 0 Set the factor 'x6' equal to zero and attempt to solve: x6 = 0 x6-0 Move all terms containing x to the left, all other terms to the right. Simplify x6 = 0 So there is NO SOLUTION Hope this help! :)

OpenStudy (anonymous):

That's not an option given

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

I just have to factor not find the solution.

OpenStudy (anonymous):

ok